What volumes of 0.1M Sodium phosphate monobasic (NaH2PO4), 0.1M sodium phosphate dibasic (Na2HPO4), and water would you have to mix to make 100mL of a .05M Phosphate buffer at pH 7.4? pKa = 7.21
Answer #1

Answer

Volume of 0.1M NaH2PO4 required = 19.6ml

Volume of 0.1M Na2HPO4 required = 30.4ml

Volume of water required = 50ml

Explanation

Henderson-Hasselbalch equation is as follows

pH = pKa + log([A-]/[HA])

A- = conjucate base , HPO42-

HA = weak acid , H2PO4-

pKa = 7.21

substituting the values

7.40 = 7.21 + log([HPO42-]/[H2PO4-])

log([HPO42-]/[H2PO4-]) = 0.19

[HPO42-]/[H2PO4-] = 1.549

[HPO42-] = 1.549× [H2PO4-]

[H2PO4-] + (1.549× [H2PO4-]) = 0.05M

2.549[H2PO4-] = 0.05M

[H2PO4-] = 0.05M/2.549

[H2PO4-] = 0.0196M

[ HPO42-] = 0.050M - 0.0196M = 0.0304M

moles of HPO42- required = (0.0304mol/1000ml)×100ml = 0.00304mol

moles of H2PO4- required = (0.0196mol/1000ml)×100ml = 0.00196mol

Volume of 0.1M NaH2PO4 required = (1000ml/0.1mol)×0.00196mol = 19.6ml

Volume of 0.1M Na2HPO4 required = (1000ml/0.10mol) × 0.00304mol = 30.4ml

Volume of water required = 100ml - (19.6ml + 30.4ml) = 50ml